OpenStudy (anonymous):

ignacio earns 5 percent simple interest annually on his savings account. How much interest will he earn each month on $1,250? Simple interest=p*r*t

OpenStudy (lgbasallote):

first..solve for the annual interest i = p*r*t where: p = 1250 t = 1 r = 0.05 can you try to solve for i?

OpenStudy (anonymous):

It will be\[ \ \huge $1,250 * (.05 ÷12) = $5.21 \]

OpenStudy (anonymous):

Get that

OpenStudy (anonymous):

yea thanks